A skating rink is designed in the shape of a rectangle with a semicircle on each end. The lenght of the rectangular portion of the rink is 44 feet and the width of the rectangular portion of the rink is 20 feet.
a.Find the perimeter of the skating rink. Round answer to the nearesth tenth.
128ft + 31.4ft(2)= 190.8ft
b. Find the total area of the skating rink. Round answer to the nearesth tenth.

Answer by scott8148(5880)

b. (44 * 20) + 100 * pi = 880 + 314.2 = 1194.2 ft^2